IN 10 years time a father will be twice as old as his daughter. 10 years he was six times as old as his daughter. How old is each now?
Answer by mananth(16949)

let fathers age now be x
daughters age now be y
10 years hence
(x+10)= 2(y+10)
x+10 = 2y+20
x-2y=10
10 years ago
(x-10)= 6(y-10)
x-10 = 6y-60
x-6y=-50
x-2y=10
x-6y=-50
subtract the equations
4y=60
y=15 daughter's age now
substitute y in x-2y=10
so x=40 present father's age